Grade 11Mechanics

If the velocity of a partial is v = At + Bt², where A and B are constants,then the distance travelled by it between 1s and 2 s is

Harshit Raj
9 Years agoGrade 11

1 Answer

Ritha
9 Years ago
when t=1 then v=A+B and when t=2 then v=2A+4B
Now acceleration(a)=(2A+4B-A-B)*1/(2-1)=A+3B
we know that v^2-u^2=2as
                  =>s=v^2-u^2*1/2a  =>s=3A^2+14AB+15B^2*1/2A+6B